1 + 5 = 12,
2 + 10 = 24,
3 + 15 = 36,
5 + 25 = 30,
then
1 + 5 = 12, false
2 + 10 = 24, false
3 + 15 = 36, false
5 + 25 = 30, true

Found 2 solutions by checkley79, vickry0804:
Answer by checkley79(3341)

1 + 5 = 12, 2 + 10 = 24, 3 + 15 = 36, 5 + 25 =60 ANS.
(1+5)2=12, (2+10)2=24, (3+15)2=36, (5+25)2=60

Answer by vickry0804(1)


1 + 5 = 12,
2 + 10 = 24,
3 + 15 = 36,
5 + 25 = 30,
then
1 + 5 = 12, false
2 + 10 = 24, false
3 + 15 = 36, false
5 + 25 = 30, true